# WarmServe: Enabling One-for-Many GPU Prewarming for Multi-LLM Serving

### v1

@@ Abstract
Deploying multiple models within shared GPU clusters is promising for improving resource efficiency in large language model (LLM) serving. Existing multi-LLM serving systems optimize GPU utilization at the cost of worse inference performance, especially time-to-first-token (TTFT). We identify the root cause of such compromise as their unawareness of future workload characteristics. In contrast, recent analysis on real-world traces has shown the high periodicity and long-term predictability of LLM serving workloads.
We propose universal GPU workers to enable one-for-many GPU prewarming that loads models with knowledge of future workloads. Based on universal GPU workers, we design and build WarmServe, a multi-LLM serving system that (1) mitigates cluster-wide prewarming interference by adopting an evict-aware model placement strategy, (2) prepares universal GPU workers in advance by proactive prewarming, and (3) manages GPU memory with a zero-overhead memory switching mechanism. Evaluation under real-world datasets shows that WarmServe improves TTFT by up to 50.8 compared to the state-of-the-art autoscaling-based system, while being capable of serving up to 2.5 more requests compared to the GPU-sharing system.
@@ 1 Introduction
The rapid evolution of large language models (LLMs) has produced various models specialized for tasks including chat [gpt4, gpt-4o, deepseek-v3, llama3], coding [claude, qwen-coder], and reasoning [deepseek-r1, openai-o3, gemini, qwen3]. This proliferation introduces significant challenges for model serving, as platforms must now concurrently host multiple models to meet varied user demands. For example, OpenAI offers more than 20 model variants through its API [openai-models].
@@
Serving multiple LLMs is challenging due to the highly dynamic workloads. As LLM loads frequently change [blitzscale, muxserve, prism], a dedicated allocation strategy (i.e., a set of GPUs is dedicated to serve a particular model) leads to low GPU utilization. To solve this problem, existing solutions can be divided into two categories: (1) autoscaling approaches that automatically scale the number of instances for each model according to current loads [serverlessllm, blitzscale, lambdascale, paraserve, atc25_torpor, atc25_deepserve], and (2) GPU sharing approaches that colocate models on the same GPUs and perform spatial and temporal sharing [li2023alpaserve, muxserve, prism, qlm].
Unfortunately, although these approaches can improve cluster-wide GPU utilization, they have non-negligible impact on inference performance. Autoscaling approaches incur waiting periods for a burst of requests, as model instances should be initialized on-demand. GPU sharing approaches have no initialization overhead, but limit the KV cache space of colocated models, which is essential for LLM serving.
@@
We attribute the inefficiency of existing approaches to the unawareness of future workload characteristics. Due to this lack of foresight, autoscaling can only be triggered after bursty requests arrive, and the model colocation strategy must be stable over time. However, real-world production traces have shown that while the short-term request arrival pattern is unpredictable, the long-term statistical characteristics of LLM requests tend to be more predictable [burstgpt, dynamollm, servegen]. Our analysis (§2.2) also confirms that LLM request demands in 5-minute windows can be effectively predicted with an accuracy of approximately 93%.
Therefore, one natural solution for efficiently serving multiple LLMs is to provision sufficient instances for models to meet the predicted maximum request loads. However, constantly preparing for peak demands leads to significant GPU consumption and low GPU utilization, as it reserves far more instances than the current load requires.
Our key insight is that we can satisfy future demand without launching the entire instance. Instead, GPUs that already contain parameters for a given model are able to provide potential serving capacity. Upon load spikes, as model weights are already resident on these GPUs, the cluster can quickly create new instances on top of them and begin inference.
This approach creates opportunities to pack backup instances for different models onto the same GPUs. A GPU can store (i.e., prewarm) parameters for multiple models because, in modern LLM deployments, model weights occupy only a small fraction of the total GPU memory. This is due to two main factors: (1) the substantial memory required by the KV cache to serve long contexts [loongserve, shuffleinfer, kunserve, deepspeed-fastgen] and facilitate prefix caching [atc25-wang, mooncake, tokenlake], and (2) the use of larger-scale parallelism to reduce latency. For example, in the official DeepSeek-V3 deployment, model weights account for 30% and 7% of GPU memory in prefill and decode instances, respectively [deepseek-v3]. Furthermore, prewarming the first several layers is sufficient for rapid startup, as the loading of subsequent layers can be overlapped with the forward computation of the loaded layers.
We propose universal GPU workers—GPU workers capable of seamlessly transforming into serving instances for any model. Specifically, we create universal GPU workers on top of idle GPUs by proactively loading weights from different models. During load spikes, these workers can be quickly reconfigured into dedicated GPU workers for specific models by evicting the weights of other models, enabling rapid startup. This approach allows a single GPU to serve as a prewarmed backup for multiple models, significantly reducing the total number of GPUs that need to be reserved.
The goal of universal GPU workers is to realize one-for-many prewarming, which has been extensively adopted in reducing cold start latency for serverless functions [sosp23_xfaas, rainbowcake, catalyzer, mitosis]. For example, one can prepare multiple universal containers so that when a request arrives, it can directly use a prewarmed one after loading a few function-specific packages [faascache, rainbowcake, sosp23_xfaas]. However, the unique characteristics of LLMs introduce two challenges to achieving one-for-many prewarming.
First, as LLMs span across multiple GPUs, a prewarmed model requires the availability of all GPUs that host its weights. Therefore, the allocation of a GPU to one model invalidates other models residing on it. Although only a single weight partition of these invalidated models becomes unavailable, the entire prewarming effort for these models is wasted because a rapid startup requires all partitions simultaneously. Given the unpredictability of short-term requests, such interference among colocated models significantly complicates the prewarming decisions for universal GPU workers.
Second, as LLM workloads are dynamic, a GPU that has just been released may be quickly reallocated to other models. In production, LLM requests can increase 5 within 2 seconds [blitzscale], which leaves a short time window for prewarming new models. Due to large LLM checkpoints, most prewarming cannot complete before the GPU is allocated again, leading to limited successful prewarming.
@@
To address these challenges, we propose WarmServe, a multi-LLM serving system designed to unleash the potential of one-for-many GPU prewarming. To reduce the prewarming interference among LLMs, WarmServe adopts an evict-aware model placement strategy that prohibits partially overlapping GPU sets during model placement. The strategy minimizes interference by tightly coupling every two colocated models, and maximizes the prewarming performance under potential evictions by isolating high-priority prewarming models.
To ensure a successful prewarming before the next GPU allocation, WarmServe proactively prewarms models before a GPU is released. Specifically, we observe that GPUs that are about to be released usually have idle memory due to fewer ongoing requests than maximum capacity. Thus, WarmServe prewarms models into those GPUs in advance, and stores the new parameters in their unused KV cache space. These GPUs can quickly turn into universal GPU workers after they are released. Since the GPU memory may simultaneously contain tensors from the served model, KV cache, and prewarming models, we propose a zero-overhead memory switching mechanism to efficiently manage these tensors, enabling GPUs to be seamlessly transformed among different types.
@@
Experiments on real-world datasets show that WarmServe reduces the tail TTFT by up to 50.8 compared to the state-of-the-art autoscaling-based system, and is able to handle up to 2.5 more requests compared to the GPU-sharing system. Furthermore, experiments show that WarmServe effectively maintains the decoding performance under various scenarios.
@@
We identify the potential of model prewarming in multi-LLM serving and propose universal GPU workers to enable one-for-many GPU prewarming.
We propose an evict-aware model placement strategy, a proactive prewarming policy, and a zero-overhead memory switching mechanism to unleash the potential of universal GPU workers.
We evaluate WarmServe comprehensively to show its effectiveness compared to state-of-the-art solutions.
@@ 2 Background and Motivation
In this section, we introduce LLM serving, analyze the long-term predictability of real-world LLM serving traces, and summarize challenges.
@@ 2.1 LLM Serving
LLM inference. LLM inference consists of two stages. The prefill stage generates the first token from the initial prompt. During this stage, the key and value vectors for each token in the prompt are computed and stored in GPUs, known as the KV cache. The decoding stage then generates the remaining outputs autoregressively. In each decoding step, the model computes and stores the key and value vectors of the last token in the sequence, and generates a new token. As LLM checkpoints are large, they are typically deployed across multiple GPUs using model parallelism, which partitions the weights and synchronizes results at specific stages during inference.
To improve GPU utilization, modern serving systems often process multiple requests in a batch. A maximum batch size is typically configured to prevent performance degradation when too many requests are processed.
@@
Multi-LLM serving. Large-scale LLM service providers, such as OpenAI [openai-models], Google [gemini], and Anthropic [anthropic], offer a catalog of multiple models tailored to different use cases. Consequently, the serving cluster must be capable of concurrently hosting these models and efficiently managing dynamic workloads for each. In this case, a static GPU allocation strategy leads to resource wastage. To address this inefficiency, a variety of multi-LLM serving systems have emerged, which fall into two categories.
@@
Autoscaling solutions dynamically control the number of instances for each model based on current loads. Upon a load spike, the system creates new instances to serve arriving requests. As instance creation is on the critical path of request handling, these systems reduce this latency by caching models locally [serverlessllm, atc25_torpor, atc25_deepserve], fetching models from peers [blitzscale, lambdascale], and distributing models across servers [paraserve]. Despite these efforts, bursty requests still suffer from long waiting latency in these systems because instances are created on demand.
@@
GPU sharing solutions share GPUs among models [li2023alpaserve, muxserve, prism, qlm]. Models are colocated on the same GPUs, and are allocated specific ratios of GPU computational power based on the traffic. In these systems, GPU sharing limits the KV cache space of each model. Additionally, these systems usually enlarge the parallelism strategy of LLMs to place more models together, degrading inference performance.
@@
WarmServe uses the autoscaling approach, i.e., it assigns instances with dedicated GPUs and scales the number of instances for each model. However, unlike existing autoscaling systems that passively create instances after bursty requests arrive, WarmServe prewarms models in advance by loading their parameters into shared idle GPUs. This synthesis of on-demand autoscaling and proactive GPU sharing allows WarmServe to harness the key benefits of each design.
Autoscaler. An autoscaler is responsible for dynamically adjusting the number of instances assigned to each model in multi-LLM serving systems [serverlessllm, blitzscale, lambdascale, paraserve, atc25_torpor, atc25_deepserve]. Typically, it periodically checks the current load and number of active instances for each model. If the resource utilization of a model falls below a predefined threshold, the autoscaler attempts to shut down some of its instances. However, since requests are often distributed in a load-balanced manner (e.g., using round-robin), most instances still have active requests and cannot be terminated immediately. To address this, the autoscaler first stops scheduling new requests to the targeted instances marked for termination, and then waits for the completion of ongoing requests. This waiting period is referred to as the grace period for an instance.
@@ 2.2 Workload Predictability
While the short-term burstiness of LLM requests is often regarded as unpredictable [li2023alpaserve, weng2022mlaas], recent analysis of real-world LLM services has revealed that the long-term statistical characteristics of requests are relatively periodic and predictable [burstgpt, dynamollm, servegen]. In this paper, we focus on the characteristics of the average and peak load, which refers to the average and maximum number of concurrent requests within a given time window, respectively.
Take peak load as an example. Figure 1 shows the values of load peaks in 5-minute windows under varying inference speeds, using data collected from the AzureConv [dynamollm] trace. The running time of a request is computed based on its input and output lengths. The results demonstrate that the load peaks follow a periodic pattern across all speeds, indicating that peak values within a given time window can be effectively predicted using historical data.
@@
To validate this assumption, we develop a simple corrective seasonal predictor (§5.1) to predict the peak load for each time window. Figure 2 shows the prediction performance of our predictor under the most dynamic load (8 Speed). By using the recorded peak loads of both previous days and recent times, the predictor achieves an average accuracy of 92.7%, which is sufficient for prewarming models. Notably, predicting the average load is even easier, with an average accuracy of 94.7%.
@@
This paper leverages the workload predictability to prewarm appropriate instances for different LLMs. By ensuring that sufficient instances are prewarmed for each LLM, most incoming requests can be served without delay. Additionally, we propose universal GPU workers to prewarm multiple LLMs simultaneously, significantly reducing the GPU resources consumed by prewarming.
2.3 Challenges
While prewarming multiple LLMs on several universal GPU workers is promising, unleashing its full potential presents two challenges listed below.
Cluster-wide prewarming interference. The large checkpoints of LLMs typically necessitate their deployment across multiple GPUs. Since a universal GPU worker is responsible for concurrently prewarming multiple models, these colocated models can experience prewarming interference. For example, Figure 3 illustrates a potential weight placement strategy when prewarming three LLMs using two GPUs. As the deployment of LLMs A and B requires two GPUs, their weights are partitioned and placed on both GPUs during prewarming.
When LLM C encounters a burst of requests, the system spawns a new instance on GPU 0 for LLM C, and evicts the parameters of LLMs A and B from that GPU. This eviction renders the first weight partition of LLMs A and B unavailable, which also invalidates their second partitions since prewarming requires all parameters to reside on GPUs. Consequently, prewarming contention arises between every two interleaved models: a prewarming hit for one model invalidates all partitions of another, even if they only share a single GPU. Such cluster-wide prewarming interference significantly complicates the design of effective model placement strategies.
Frequent GPU reallocation. LLM workloads are highly dynamic, with request rates capable of increasing 5 within just 2 seconds [blitzscale]. This leads to a high frequency of GPU allocation and deallocation, where idle GPUs are quickly assigned to new models. Such rapid reallocation cycles leave a narrow time window for prewarming new models. As loading models into GPUs is time-consuming due to limited PCIe bandwidth and the large size of LLM checkpoints, most prewarming attempts fail to complete before GPUs are reallocated, leading to low prewarming hit probability.
3 WarmServe Overview
To address these challenges, we design and build WarmServe, a multi-LLM serving system that strategically prewarms models to enable one-for-many prewarming. Figure 4 shows the architecture of WarmServe. WarmServe employs a global manager to control a pool of GPUs (i.e., GPU workers), categorized as idle, universal, or dedicated. The global manager dynamically loads or evicts model weights on these GPUs, adapting their roles in different scenarios. Specifically, idle GPU workers transition into universal workers after prewarming several LLMs. Universal workers, in turn, become dedicated GPU workers when a new instance of a prewarmed model is launched. Once the instance is released, these workers revert to idle status and begin prewarming new models.
WarmServe adopts an evict-aware model placement strategy (§5.2) to mitigate prewarming interference. This strategy restricts GPU sharing to occur only when the GPU set of one prewarmed LLM fully contains that of another, ensuring the models are tightly coupled. Additionally, WarmServe computes the scores for the prewarming replicas of each model based on load predictions (§5.1). It then isolates high-score replicas to enhance the overall prewarming effectiveness.
To address the challenge of frequent GPU reallocation, WarmServe proactively prewarms models on dedicated GPU workers (§4.1). This strategy is applied to grace-period instances, and the weights of new models are stored in unused KV cache space. Once the instance is terminated, its dedicated GPU workers can seamlessly transform into universal workers without additional weight loading. To ensure the flexible coexistence of serving model parameters, prewarmed model weights, and KV cache, WarmServe dynamically switches the virtual addresses of GPU memory and adjusts address mappings during prewarming and allocation (§4.2). This memory management technique hides all time-consuming operations behind other actions to achieve zero overhead.

### v2

@@ Abstract
Deploying multiple models within shared GPU clusters is a key strategy to improve resource efficiency in large language model (LLM) serving. Existing multi-LLM serving systems improve GPU utilization at the cost of degraded inference performance, particularly time-to-first-token (TTFT). We attribute this degradation to the lack of awareness regarding future workload characteristics. In contrast, recent analyses have shown the strong periodicity and long-term predictability of real-world LLM serving workloads. In this paper, we propose one-for-many GPU prewarming, which proactively loads parameters from multiple models onto GPUs based on workload forecasts. These prewarmed weights enable the system to promptly instantiate serving instances upon encountering request bursts. We design and implement WarmServe, a multi-LLM serving system incorporating three key techniques: (1) a model placement algorithm that optimizes prewarming decisions to minimize cross-model prewarming interference, (2) a KV cache reservation strategy that repurposes idle KV cache space on running GPUs for prewarming new models, and (3) an efficient GPU memory switching mechanism for tensor management. Evaluation on real-world datasets shows that WarmServe reduces tail TTFT by up to 50.8 compared to the state-of-the-art autoscaling-based system, while supporting up to 2.5 higher request throughput than the GPU-sharing system.
@@ 1 Introduction
The rapid evolution of large language models (LLMs) has led to a diverse ecosystem of specialized models tailored for tasks such as chat (OpenAI, 2023a, 2025a; DeepSeek-AI, 2025b; Grattafiori et al., 2024), coding (Anthropic, 2025a; Team, 2025), and reasoning (DeepSeek-AI, 2025a; OpenAI, 2023b; Google, 2025). This proliferation introduces significant challenges for model serving, as platforms must now concurrently host a multitude of model variants to meet varied user demands. For example, OpenAI offers more than 20 model variants through its API (OpenAI, 2025b).
@@
Efficiently serving multiple LLMs is challenging due to highly dynamic workloads. Since request volumes fluctuate frequently (Zhang et al., 2025; Duan et al., 2024; Yu et al., 2025c), a dedicated allocation strategy (i.e., a set of GPUs is dedicated to serving a particular model) leads to severe GPU under-utilization during idle periods. To solve this problem, existing solutions can be divided into two categories: (1) autoscaling approaches that scale the number of instances for each model according to current loads (Fu et al., 2024; Zhang et al., 2025; Yu et al., 2025b; Lou et al., 2026; Yu et al., 2025a; Hu et al., 2025; Zhu et al., 2025; Liu et al., 2025; Mei et al., 2026), and (2) GPU sharing approaches that colocate multiple models on the same GPU through spatial and temporal sharing (Li et al., 2023; Duan et al., 2024; Yu et al., 2025c; Patke et al., 2024; Gao et al., 2025; Xiang et al., 2025).
Unfortunately, although these approaches can improve cluster-wide GPU utilization, they impose a non-negligible impact on inference performance. Autoscaling suffers from significant cold-start latency, as initializing new instances on-demand during request bursts is time-consuming. Conversely, GPU sharing avoids initialization delays but severely constricts the KV cache capacity available to each model—a critical resource for maintaining high throughput and handling long sequences in LLM serving.
@@
We attribute the limitations of existing approaches to the lack of awareness of future workload characteristics. Due to this lack of foresight, autoscaling is only triggered after bursty requests arrive, and the model colocation strategy must be stable over time. However, recent production traces reveal a key insight: while short-term request arrivals are inherently stochastic, the long-term statistical trends of LLM workloads exhibit strong periodicity (Wang et al., 2025; Stojkovic et al., 2025; Xiang et al., 2026). Our analysis (Section 2.2) confirms this, demonstrating that peak LLM request volumes within 5-minute windows can be predicted with an average relative error of 7.3%.
Leveraging this high predictability, we propose a proactive autoscaling approach centered on LLM prewarming to bridge the gap between resource efficiency and inference performance. Instead of reacting to request arrivals, the system provisions model replicas in anticipation of forecasted load spikes. When the predicted future demand of a specific model exceeds its current capacity, the system proactively launches backup replicas on idle GPUs, enabling the cluster to directly serve potential bursts without delay.
Nevertheless, maintaining exclusive backup replicas for every model is resource-intensive and restricts prewarming scalability. To address this, we introduce one-for-many GPU prewarming, which loads parameters from multiple models into the memory of a single GPU. Once a specific model encounters a request burst, it can immediately instantiate an active serving instance using its pre-loaded weights. The system then evicts non-target parameters to guarantee exclusive GPU access for the active instance. Since evicting irrelevant weights is substantially faster than loading weights on demand, one-for-many GPU prewarming improves inference performance while reducing the GPU resources required for prewarming.
While prewarming has been extensively adopted to reduce cold start latency for serverless functions (Sahraei et al., 2023; Yu et al., 2024; Du et al., 2020; Wei et al., 2023), adapting this technique to LLMs introduces two unique challenges:
1.
Cross-model prewarming interference. LLMs are often distributed across multiple GPUs (e.g., via tensor parallelism). A prewarmed model is only functional if all its constituent GPUs are available. Consequently, allocating a single GPU to one model evicts and invalidates multiple other prewarmed models that reside on this GPU, leading to cross-model prewarming interference that complicates placement decisions.
2.
Transient prewarming windows. LLM workloads are extremely volatile. For example, LLM requests can surge 5 within two seconds in production (Zhang et al., 2025). This leaves a narrow window to prepare new models. Given the massive size of LLM checkpoints, most prewarming attempts fail to complete before the next allocation cycle, resulting in wasted effort and failed prewarming.
@@
To address these challenges, we present WarmServe, a multi-LLM serving system designed to unleash the potential of one-for-many GPU prewarming. To mitigate cross-model interference, WarmServe formulates the model placement problem and employs a prewarming algorithm that maximizes the prewarming efficacy under potential evictions by strategically isolating high-priority models.
To guarantee that prewarming succeeds within transient windows, WarmServe proactively initiates prewarming before a GPU is officially released. We observe that GPUs approaching release typically have idle memory as their active load falls below peak capacity. Leveraging this slack, WarmServe loads new parameters into the unused KV cache space of these still-active GPUs. Once released, these GPUs can quickly transition into a serving instance of the pre-loaded models, regardless of when reallocation occurs. We develop a strategy to compute the required KV cache for ongoing requests on these GPUs, and introduce a GPU memory switching mechanism to efficiently manage diverse tensors.
@@
Experiments on real-world datasets show that WarmServe reduces the tail TTFT by up to 50.8 compared to the state-of-the-art autoscaling-based system, and is capable of handling up to 2.5 more requests compared to the GPU-sharing system.
@@
We identify the potential of model prewarming in multi-LLM serving and introduce one-for-many GPU prewarming to efficiently prepare backup model replicas based on future workload forecasts.
We formulate the joint problem of model placement and proactive prewarming, developing an optimized placement algorithm alongside a dynamic KV cache reservation strategy. We also propose a GPU memory switching mechanism for efficient tensor management.
We conduct a comprehensive evaluation of WarmServe, demonstrating its superior performance and efficiency compared to state-of-the-art solutions.
@@ 2 Background and Motivation
In this section, we introduce LLM serving and analyze the long-term predictability of real-world LLM serving traces.
@@ 2.1 LLM Serving
LLM Inference. LLM inference consists of two stages. The prefill stage generates the first token from the initial prompt. During this stage, the key and value vectors for each token in the prompt are computed and stored in GPUs, known as the KV cache. The decoding stage then generates the remaining outputs autoregressively. In each decoding step, the model computes and stores the key and value vectors of the last token in the sequence, and generates a new token. As LLM checkpoints are large, they are typically deployed across multiple GPUs using model parallelism, which partitions the weights and synchronizes results at specific stages during inference. To improve GPU utilization, modern serving systems process multiple requests in a batch. A maximum batch size is typically configured to prevent performance degradation when too many requests are processed.
@@
Multi-LLM Serving. Large-scale LLM service providers, such as OpenAI (OpenAI, 2025b), Google (Google, 2025), and Anthropic (Anthropic, 2025b), offer a catalog of multiple models tailored to different use cases. Consequently, the serving cluster must be capable of concurrently hosting these models and efficiently managing dynamic workloads for each. In this case, a static GPU allocation strategy leads to significant under-utilization of GPUs since LLMs have fluctuating real-time loads. To address this inefficiency, a variety of multi-LLM serving systems have emerged, which fall into two categories:
@@
Autoscaling solutions that dynamically control the number of instances for each model based on current loads. Upon a load spike, the system creates new instances to serve arriving requests. As instance creation is on the critical path of request handling, these systems reduce creation latency by caching models locally (Fu et al., 2024; Yu et al., 2025a; Hu et al., 2025; Zhu et al., 2025; Liu et al., 2025), fetching models from peers (Zhang et al., 2025; Yu et al., 2025b), and distributing models across servers (Lou et al., 2026). Despite these efforts, bursty requests still suffer from long waiting latency in these systems because instances are created on demand.
@@
GPU sharing solutions that share GPUs across models (Li et al., 2023; Duan et al., 2024; Yu et al., 2025c; Patke et al., 2024; Gao et al., 2025; Xiang et al., 2025). Models are initially deployed and colocated on the same GPUs, and are allocated specific ratios of GPU computational power based on the traffic. In these systems, GPU sharing limits the KV cache space of each model. Additionally, these systems usually increase the degree of parallelism for LLMs to colocate more models, which degrades inference performance.
@@
WarmServe adopts the autoscaling approach, i.e., it assigns dedicated GPUs to instances and scales the instances for each model. However, unlike existing autoscaling systems that passively create instances after bursty requests arrive, WarmServe prewarms models in advance by loading their parameters into shared idle GPUs. This synthesis of on-demand autoscaling and shared GPU preparation allows WarmServe to harness the key benefits of each design.
@@ 2.2 Workload Predictability
While the short-term burstiness of LLM requests is often regarded as unpredictable (Li et al., 2023; Weng et al., 2022), recent analyses of real-world LLM services have revealed that the long-term statistical characteristics of requests are relatively periodic and predictable (Wang et al., 2025; Stojkovic et al., 2025; Xiang et al., 2026). In this paper, we focus on the characteristics of the average and peak loads, which refer to the average and maximum number of concurrent requests within a specific time window, respectively.
We validate workload predictability by taking peak load as a representative case. Our analysis leverages production traces including AzureConv (Stojkovic et al., 2025), AzureCode (Stojkovic et al., 2025), BurstGPT (Wang et al., 2025), and ServeGen111We choose the m-small workload as it is the most popular model. (Xiang et al., 2026). Figure 1 shows the values of load peaks in 5-minute windows for different traces. The results demonstrate that the load peaks of a single LLM follow a periodic pattern in all traces, indicating that peak values within a given time window can be effectively predicted using historical data.
@@
We further develop a simple corrective seasonal predictor to predict the peak load for each time window (Section 3.1). Figure 2 shows the prediction performance of our predictor on AzureConv. By using the recorded peak loads of both previous days and recent windows, the predictor achieves an average accuracy of 92.7%, which is sufficient for model prewarming. Notably, predicting the average load is even easier, with an average accuracy of 94.7%. The detailed results for all traces are shown in Section 4.6.
@@
This paper leverages the per-model workload predictability to prewarm appropriate instances for different LLMs. By ensuring that sufficient instances are prewarmed for each LLM, most incoming requests can be served without delay. Additionally, we propose one-for-many GPU prewarming to prepare multiple LLMs simultaneously, significantly reducing the GPU resources consumed by prewarming.
3 Method
We design and build WarmServe, a multi-LLM serving system that performs one-for-many GPU prewarming based on workload predictions. Figure 3 shows the architecture of WarmServe. WarmServe classifies GPU workers into three types: idle, universal, and dedicated. For idle GPU workers, WarmServe prewarms several LLMs on them by loading model weights onto GPU memory, converting these workers into universal GPU workers. Prewarming a model often requires multiple GPUs, each holding part of the weights. When one of the prewarmed models encounters request bursts, the loaded model weights enable the GPUs to quickly create engines and start serving, becoming dedicated GPU workers. Prewarmed weights of other models on these GPUs are evicted. WarmServe also allows prewarming to happen on dedicated GPU workers by storing weights in their unused KV cache space.
At runtime, WarmServe predicts the future workload for each LLM (Section 3.1), and leverages a model placement algorithm (Section 3.2) to generate prewarming actions. For dedicated GPUs that will be released soon due to decreased loads, WarmServe proactively prewarms new models onto their unused KV cache so that the GPUs can directly become universal GPU workers after being released (Section 3.3). Finally, WarmServe adopts a GPU memory switching mechanism to manage GPU memory (Section 3.4), ensuring the flexible coexistence of serving model parameters, prewarmed model weights, and KV cache.